OpenStudy (anonymous):

What is the value of f(2) for the function f(x) = 5x + 3?

OpenStudy (anonymous):

plug the x in to the equation

OpenStudy (anonymous):

put x=2 and you will get your answer as 13

OpenStudy (anonymous):

then solve

OpenStudy (***[isuru]***):

f(x) = 5x +3 now the question asks for the f(2) So u must substitute 2 instead of x in f(x) .... f(x) contain only one term which include x - i.e "5x" now substitute 2 instead of x in "5x" and u will get a expression like this 5x +3 -----> (5*2) +3 = 13 yep! , that's the value of f(2) ..... Did u get it now ?

OpenStudy (anonymous):

kinda can you show me another one so i can understand it What is the value of g[f(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (***[isuru]***):

Let's start with f(x)... f(x) = 2x + 1 the question requests f(5) so.... substitute 5 instead of x in f(x) 2x + 1 -----> (2*5) +1 which is 11 so f(5) = 11 Now let's move to the second part g(x) = 2x - 5 now the question asks for g[f(5)] now substitute f(5) instead of x in g(x) 2x - 5 ------------> [2*f(5)] -5 since f(5) = 11 [2*f(5)] -5 = (2*11) -5 = 17 yep that's it! g[f(5)] = 17